OpenStudy (anonymous):

Please help! I need to create 3 Literal Equation with 3 different objects. A baseball, a basketball, and a football. All variables in the formula must be in the correct position and the Final solution must be correct. XOxo

OpenStudy (anonymous):

@GreenCat @ProLearner @paki @sleepyjess I am not sure where to begin or how to creatively make literal equations out of these items

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I think "literal equation" is some new goofy terminology. Do you have any background or info on what they mean ? Anything ?

OpenStudy (phi):

I can imagine writing down the formulas for the volume of each object. But I can't make sense of "solving" them.

OpenStudy (texaschic101):

I found this.. This process of solving a formula for a given variable is called "solving literal equations". One of the dictionary definitions of "literal" is "related to or being comprised of letters", and variables are sometimes referred to as literals. So "solving literal equations" seems to be another way of saying "taking an equation with lots of letters, and solving for one letter in particular." phi...does this mean we can make a 3 variable system of equations and solve by elimination ?

OpenStudy (anonymous):

@phi thank you for your time. I was given an example of a literal equation and the entire equation was made up of variables. I had to reverse the order of operations to solve for the desired variable. Could you give me a formula for the volume of one of the objects please?

OpenStudy (anonymous):

@texaschic101 thank you for your time as well.

OpenStudy (phi):

@texaschic101 that is helpful, but I still don't understand the question. I think this might be "common core" ??

OpenStudy (phi):

both the baseball and basketball are spheres, so \[ V = \frac{4}{3} \pi r^3 \] where r is the radius the football is a "prolate spheriod" with volume \[ V = \frac{4}{3} \pi a^2 c \] where a is the radius of the short way round = 3.3 inches and c is the radius of the long way round is about 4.5 inches